OpenStudy (anonymous):

if you subtract 8 from me then subtract 2 more, you get 5

OpenStudy (akashdeepdeb):

Okay let us say I am a number x [Like Chemical X] So if you subtract 8 from me that is x-8 And then subtract 2 from me again x -8 -2 you get 5 So writing this like Math --> x -8-2 = 5 x-10=5 x - 10 + 10 = 5 + 10 [ADD 10 TO BOTH SIDES] x = 15 Yayy! So I am 15! You subtract 8 from me and get 7! You subtract 2 more you get 5 [YIPPEE!] Understood? :)

OpenStudy (anonymous):

u smart

OpenStudy (akashdeepdeb):

@RamyaR Did you get it? @anajisselle Not quite! ,But thanks! :D